OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!! If f(x)=2x^2-4x-3 and g(x)=(x-4)(x+4),find f(x)-g(x) Answers: 4x+13 x^2-4x+13 3X^2-4x+13 -4x^2-4x+13

OpenStudy (anonymous):

@studydood66

OpenStudy (anonymous):

he can probably halp

OpenStudy (anonymous):

74895308

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Thank you!!!! @studydood66

OpenStudy (anonymous):

@studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66

OpenStudy (anonymous):

Could you help me with other questions? @studydood66

OpenStudy (anonymous):

i think its right but im not positive

OpenStudy (anonymous):

and yea

OpenStudy (anonymous):

@studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66

OpenStudy (anonymous):

@studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66 @studydood66

OpenStudy (anonymous):

lol srry

OpenStudy (anonymous):

If f(x)=(x+1)^-1 and g(x)=x-2, what is the domain of f(x)/g(x)

OpenStudy (anonymous):

that im not shure

OpenStudy (anonymous):

lol you good @megax

OpenStudy (anonymous):

will posting the answers help? @studydood66

OpenStudy (anonymous):

im looking it up

OpenStudy (anonymous):

f(x)/ g(x) = (1/x )/ ((x+1)/(x-2))

OpenStudy (anonymous):

all values of x (infinity,-1),(-12), and (2,infinity) (-infinity, 2) and (2,infinity) (-infinity,-1] and [2,infinity) @studydood66

OpenStudy (anonymous):

yea thats the answer

OpenStudy (anonymous):

its not on there

OpenStudy (anonymous):

ok im not interaly shure about it what grade is that

OpenStudy (anonymous):

12 grade lol @studydood66

OpenStudy (anonymous):

thats why i have to look it up im 7th

OpenStudy (anonymous):

oh dang i feel so stupid XD @studydood66 but thank you for the help

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can you try helping me with another math question? @studydood66